# Non-linear degenerate parabolic flow equations and a finer differential structure on Wasserstein spaces

### v1

@@ 3.6.1 Average Flow Approximation Series (AFAS)
For $(M,g)=(\mathbb{R}^{n},g^{eucl})$, both statements follow by induction using (3.56) in Lemma 3.31. Let $t\in[0,T]$. For the first statement, we start with $\mu=\nu=\mu_{t}$ and in every step, we set $v_{1}=\mu$ and $V[v_{2}]=\delta_{0}$. For the second statement, we start with $\mu=\mu_{t}$ and $\nu=\delta_{0}$ and in every step, we set $v_{1}=\mu$ and $V[v_{2}]=\delta_{0}$. Note that for all $\mu\in\mathcal{W}_{p}(M)$,
$\mathcal{W}_{p}(V[\mu],\delta_{0})$ $\leq R,$ (3.78)
$\mathcal{W}_{p}(\mu,\delta_{0})^{p}$ $=\mathcal{M}_{p}(\mu).$
Finally, we obtain the estimate for an arbitrary manifold with bounded geometry by a straightforward application of Nash’s embedding theorem. $\square$
Finally, Lemma 3.31 allows us to give uniform upper bounds for the difference of the pp–th and the qq–th moments of an AFAS.
@@
Let $T>0$, $p\geq q\geq 2$, $\mu_{0}\in\mathcal{W}_{p}(M)$. Let $\mathcal{}\mathfrak{P}$ be a partition of $[0,T]$. Let $V$ be a vector field probability. Assume that there exists $R>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $V[\mu]$ is supported in the centered ball with radius $R$. Then, there exists $\delta>0$ and a constant $\mathcal{K}>0$ depending only on $T,p,R$ and $M$, such that, if $\mathfrak{p}(\mathcal{}\mathfrak{P})\leq\delta$, for all $t\in[0,T]$,
$\biggl{|}\mathcal{M}_{p}(\mu^{\mathcal{}\mathfrak{P}}_{t})-\mathcal{M}_{q}(\mu^{\mathcal{}\mathfrak{P}}_{t})\biggr{|}\leq\biggl{|}\mathcal{M}_{p}(\mu_{0})-\mathcal{M}_{q}(\mu_{0})\biggr{|}+C(p-q)t.$ (3.79)
@@
The result follows from a computation similar to the one in the proof of Lemma 3.31. Indeed,
@@
$\qquad+\bigl{(}p\|x-x_{0}\|^{p-2}-q\|x-x_{0}\|^{q-2}\bigr{)}\bigl{<}x-x_{0},X(x)\bigr{>}t$
$\qquad+\int_{0}^{t}(t-s)\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}\biggl{(}t\mapsto\bigl{\|}\rho^{X}_{t}(x)-x_{0}\bigr{\|}^{p}-\bigl{\|}\rho^{X}_{t}(x)-x_{0}\bigr{\|}^{q}\biggr{)}(s)\operatorname{d}s,$
@@
$\qquad=\biggl{(}p(p-2)\bigl{\|}\rho^{X}_{s}(x)-x_{0}\bigr{\|}^{p-2}-q(q-2)\bigl{\|}\rho^{X}_{s}(x)-x_{0}\bigr{\|}^{q-2}\biggr{)}\biggl{<}\frac{\rho^{X}_{s}(x)-x_{0}}{\|\rho^{X}_{s}(x)-x_{0}\|},$
$\qquad\frac{\operatorname{d}}{\operatorname{d}t}\biggl{(}t\mapsto\rho^{X}_{t}(x)\biggr{)}(s)\biggr{>}^{2}$
$\qquad+\biggl{(}p\bigl{\|}\rho^{X}_{s}(x)-x_{0}\bigr{\|}^{p-2}-q\bigl{\|}\rho^{X}_{s}(x)-x_{0}\bigr{\|}^{q-2}\biggr{)}\biggl{[}\biggl{\|}\frac{\operatorname{d}}{\operatorname{d}t}\biggl{(}t\mapsto\rho^{X}_{t}(x)\biggr{)}(s)\biggr{\|}^{2}$
$\qquad+\biggl{<}\rho^{X}_{s}(x)-x_{0},\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}\biggl{(}t\mapsto\rho^{X}_{t}(x)\biggr{)}(s)\biggr{>}\biggr{]}.$
@@
$\qquad\leq(p-q)\biggl{|}\log\bigl{(}\bigl{\|}\rho^{X}_{s}(x)-x_{0}\bigr{\|}\bigr{)}\biggr{|}\max\biggl{(}\bigl{\|}\rho^{X}_{s}(x)-x_{0}\bigr{\|}^{p-2},\bigl{\|}\rho^{X}_{s}(x)-x_{0}\bigr{\|}^{q-2}\biggr{)}.$
@@
$\biggl{|}\log\bigl{(}\bigl{\|}\rho^{X}_{s}(x)-x_{0}\bigr{\|}\bigr{)}\biggr{|}\bigl{\|}\rho^{X}_{s}(x)-x_{0}\bigr{\|}^{q-2}\leq\bigl{\|}\rho^{X}_{s}(x)-x_{0}\bigr{\|}^{q-2+\delta}+\bigl{\|}\rho^{X}_{s}(x)-x_{0}\bigr{\|}^{q-2-\delta}.$ (3.83)
@@
$\qquad\leq\biggl{|}\mathcal{M}_{p}(\mu)-\mathcal{M}_{q}(\mu)\biggr{|}+2^{p-2}\bigl{(}8p^{2}R^{2}+6R^{2}p+3R^{2}p\bigr{)}\bigl{(}\mathcal{M}_{p}(\mu)+1\bigr{)}(p-q)t^{2}$
$\qquad+\frac{2^{2(p-1)}\bigl{(}8p^{2}R^{2}+6R^{2}p+3R^{2}p\bigr{)}R^{2}}{6}(p-q)t^{3}.$
Now, the result follows by induction on the steps in $\mathcal{}\mathfrak{P}$ using the second inequality in Corollary 3.33. $\square$
@@ 3.6.2 First Main Result: Generalised Central Limit Theorem—weak version
Let $T>0$ and $\mu_{0}\in\mathcal{W}_{p}(M)$. Let $\bigl{(}\mathcal{}\mathfrak{P}(\frac{1}{N})\bigr{)}_{N\in\mathbb{N}}$ be a family of partitions of $[0,T]$. Let $V$ be a uniformly continuous probability vector field on $\mathcal{W}_{p}(M)$ for $p>2$. Assume that $V$ has uniformly bounded pp–th moments, e.g, there exists a constant $B>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $\mathcal{M}_{p}(V[\mu])<B$.
@@
    Assume that $(M,g)$ is of bounded geometry. Assume furthermore, that there exists $R>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $V[\mu]$ is supported in the centered ball with radius $R$. Then, there exists a subsequence $\biggl{(}\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N_{k}})}\biggr{)}_{k\in\mathbb{N}}$ of the Average Flow Approximation Series $\biggl{(}\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N})}\biggr{)}_{N\in\mathbb{N}}$ with initial value $\mu_{0}$ (See Definition 3.27) which uniformly converges with respect to the pp–Wasserstein distance to a $\frac{1}{p}$–Hölder–continuous smooth solution $\mu:[0,T]\to\mathcal{W}_{p}(M)$ of (3.14) with initial value $\mu_{0}$.
@@
    Now, we only assume that there exists $q<p$ such that $V$ is continuous for the $\mathcal{W}_{q}$–distance. Then, there exists a subsequence $\biggl{(}\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N_{k}})}\biggr{)}_{k\in\mathbb{N}}$ of the Linear Average Flow Approximation Series with initial value $\mu_{0}$ which converges to a smooth solution $\mu:[0,T]\to\mathcal{W}_{p}(M)$ of (3.14) with initial value $\mu_{0}$.
@@
Let $p>2$. The proof consists of three main steps.
@@
Take $N\in\mathbb{N}$ and $l\in\llbracket 0,N-1\rrbracket$. Suppose $\frac{\delta^{t}_{N}}{2}\leq 1$. Then for all $t,s\in[l\delta^{t}_{N},\frac{2l+1}{2}\delta^{t}_{N}]$, $|t-s|\leq\sqrt{|t-s|}$. Note that since $V$ has pp–th moments bounded by $B>0$, it has in particular first and second moments bounded by $B+1$. Then, Lemma 3.29 yields that
$\mathcal{W}_{p}(\mu^{N}(t),\mu^{N}(s))$ $\leq\mathcal{M}_{1}(V[\mu^{N}(l\delta^{t}_{N})])\sqrt{|t-s|}$ (3.85)
$\leq(B+1)\sqrt{|t-s|}$
Furthermore, by Lemma 3.30, for all $t,s\in[\frac{2l+1}{2}\delta^{t}_{N},(l+1)\delta^{t}_{N}]$,
$\mathcal{W}_{p}(\mu^{N}(t),\mu^{N}(s))$ $\leq\sqrt{|t-s|}(\|\overline{V}[\mu^{N}(l\delta^{t}_{N})]\|_{\infty}+\mathcal{M}_{p}(V[\mu^{N}(l\delta^{t}_{N})])^{\frac{1}{p}})$ (3.86)
$\leq\sqrt{|t-s|}(3(B+1)+B^{\frac{1}{p}})$
Hence, there exists $\mathcal{B}>0$ such that, whenever $|t-s|\leq\frac{1}{2}\delta^{t}_{N}$, we have
@@
In the second and most technical step, we prove that the Average Flow Approximation Series from Definition 3.27 satisfies (3.14) in the limit.
Take $N\in\mathbb{N}$, and let $(x_{i})_{i\in\llbracket 0,I_{N}\rrbracket}\coloneqq\mathcal{}\mathfrak{P}(\frac{1}{N})$ indexed in increasing order, $s\in[0,T]$ and denote by $l^{N}(s)\in\llbracket 0,I_{N}-1\rrbracket$ the maximal integer satisfying $x_{l^{N}(s)}\leq s$ and $\varphi\in\mathcal{C}_{c}^{\infty}(M)$. To alleviate notations, we shall write $\mu^{N}\coloneqq\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N})}$. Then,
@@
$\qquad+\int_{M}\varphi\operatorname{d}(f^{V}_{x_{i+1}-x_{i}}(\mu^{N}(x_{i}))-\mu^{N}(x_{i}))$
@@
$\qquad+\int_{M}\int_{\mathcal{X}(M)}\varphi\biggl{(}\rho^{\overline{X}[\mu^{N}(x_{i})]}_{\sqrt{x_{i+1}-x_{i}}}(x)\biggr{)}-\varphi(x)\operatorname{d}V[\mu^{N}(x_{i})](X)\operatorname{d}\mu^{N}(x_{i})(x).$
Now, note that for all $X\in W^{1,\infty}(M)$, $d_{\infty}(\rho^{X}_{t},\operatorname{id})\leq t\|X\|_{\infty}$. Hence $\varphi\circ\rho^{X}_{t}$ is compactly supported. Furthermore, there exists a function $K:\mathcal{X}(M)\to\mathbb{R}$ bounded by $\overline{B}$ such that
@@
$\qquad\int_{\mathcal{X}(M)}\varphi(\rho^{\overline{X}[\mu^{N}(x_{i})]}_{\sqrt{x_{i+1}-x_{i}}}(x))-\varphi(x)\operatorname{d}V[\mu^{N}(x_{i})](X)$ (3.91)
$=\int_{\mathcal{X}(M)}\sqrt{x_{i+1}-x_{i}}\mathcal{L}_{\overline{X}[\mu^{N}(x_{i})]}\varphi(x)+\tfrac{1}{2}(x_{i+1}-x_{i})\mathcal{L}^{2}_{\overline{X}[\mu^{N}(x_{i})]}\varphi(x)$
$\qquad+K(X)\|\varphi\|_{\mathcal{C}^{3}}(x_{i+1}-x_{i})^{\frac{\min(3,p)}{2}}\|{\overline{X}[\mu^{N}(x_{i})u]}\|_{W^{2,\infty}}^{\min(3,p)}\operatorname{d}V[\mu^{N}(x_{i})](X)$
$=\int_{\mathcal{X}(M)}\sqrt{x_{i+1}-x_{i}}\mathcal{L}_{\overline{X}[\mu^{N}(x_{i})]}\varphi(x)+\tfrac{1}{2}(x_{i+1}-x_{i})\mathcal{L}^{2}_{\overline{X}[\mu^{N}(x_{i})]}\varphi(x)\operatorname{d}V[\mu^{N}(x_{i})](X)$
$\qquad+(x_{i+1}-x_{i})^{\frac{\min(3,p)}{2}}\int_{\mathcal{X}(M)}K(X)\|\varphi\|_{\mathcal{C}^{3}}\|{\overline{X}[\mu^{N}(x_{i})]}\|_{W^{2,\infty}}^{\min(3,p)}\operatorname{d}V[\mu^{N}(x_{i})](X).$
@@
$\int_{\mathcal{X}(M)}$ $\mathcal{L}_{\overline{X}[\mu^{N}(x_{i})]}\varphi(x)\operatorname{d}V[\mu^{N}(x_{i})](X)$ (3.92)
$=\int_{\mathcal{X}(M)}\operatorname{d}_{x}\varphi(\overline{X}[\mu^{N}(x_{i})](x))\operatorname{d}V[\mu^{N}(x_{i})](X)$
$=\int_{\mathcal{X}(M)}\operatorname{d}_{x}\varphi(v-\overline{V}[\mu^{N}(x_{i})](x))\operatorname{d}(\pi_{x})_{*}V[\mu^{N}(x_{i})](v)$
$=\operatorname{d}_{x}\varphi\biggl{(}\int_{\mathcal{X}(M)}v\operatorname{d}(\pi_{x})_{*}V[\mu^{N}(x_{i})](v)-\overline{V}[\mu^{N}(x_{i})](x)\biggr{)}=0,$

### v2

@@ 3.6.3 Averaged flow operators—Lipschitz measure–dependence
The statement follows by induction using (3.55) in Lemma 3.34. Let $t\in[0,T]$. We start with $\mu=\nu=\mu_{t}$ and in every step, we set $v_{1}=\mu$ and $V[v_{2}]=\delta_{0}$. Note that for all $\mu\in\mathcal{W}_{p}(M)$,
$\mathcal{W}_{p}(V[\mu],\delta_{0})$ $\leq R.$ (3.81)
$\square$
Furthermore, we obtain Lipschitz regularity for the pp–th moment of an (AFAS) in manifolds with bounded geometry.
Corollary 3.37.
Let $(M,g)$ be of bounded geometry, $T>0$, $p\geq 2$, $\mu_{0}\in\mathcal{W}_{p}(M)$. Let $\mathcal{}\mathfrak{P}$ be a partition of $[0,T]$, $V$ a probability vector–field. Assume that there exists $R>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $V[\mu]\in\mathcal{W}_{p+1}(\mathcal{X}^{2,\infty}(M))$ and $\mathcal{M}_{p+1}(V[\mu])\leq R^{p+1}$. Then:
  • The function $t\mapsto\mathcal{M}_{p}(\mu^{\mathcal{}\mathfrak{P}}_{t})$ is Lipschitz with respect to the pp–Wasserstein distance.
  • More precisely, for all $\varepsilon>0$, there exists $\delta>0$ such that, if the path–length $\mathfrak{p}(\mathcal{}\mathfrak{P})\leq\delta$, for all $t,s\in[0,T]$, we have
    $\bigl|\mathcal{M}_{p}(\mu^{\mathcal{}\mathfrak{P}}_{t})\bigr|\leq C(M)\bigl(\mathcal{M}_{p}(\mu_{0})+p2^{p-1}R\bigl(2pR+1\bigr)t\bigr)e^{\bigl(p2^{p-1}R\bigl(2pR+1\bigr)+\varepsilon\bigr)t}$ (3.82)
    $\hskip 18.49988pt+\varepsilon t,$
    and, in particular, there exists a constant $\mathcal{H}$ depending on $\mathcal{M}_{p}(\mu_{0})$, $T$, $R$, $p$ and $M$, such that
    $\bigl|\mathcal{M}_{p}(\mu^{\mathcal{}\mathfrak{P}}_{t})-\mathcal{M}_{p}(\mu^{\mathcal{}\mathfrak{P}}_{s})\bigr|\leq\mathcal{H}|t-s|,$ (3.83)
Proof.
For $(M,g)=(\mathbb{R}^{n},g^{eucl})$, we proceed as in the proof of Lemma 3.33. Let $t,R>0$. A lengthy computation shows that for all $\mu\in\mathcal{W}_{p}(M)$, $V\in\mathcal{W}_{p+1}(\mathcal{X}^{2,\infty}(M))$ with $\mathcal{M}_{p+1}(M)\leq R^{p+1}$,
$\biggl|\mathcal{M}_{p}\bigl(f^{V}_{t}(\mu)\bigr)-\mathcal{M}_{p}(\mu)\biggr|$ (3.84)
$\hskip 18.49988pt\leq p2^{p-2}\biggl((p-1)\mathcal{M}_{p-2}(\mu)\mathcal{M}_{2}(V[\mu])+\mathcal{M}_{p-1}(\mu)\mathcal{M}_{2}(V[\mu])\biggr)t^{2}$
$\hskip 18.49988pt+\frac{2^{p-2}}{p-1}\biggl((p-1)\mathcal{M}_{p}(V[\mu])+\mathcal{M}_{1}(\mu)\mathcal{M}(V[\mu])\biggr)t^{p}$
$\hskip 18.49988pt+\frac{2^{p-2}}{p+1}\mathcal{M}_{p+1}(V[\mu])s^{p+1}+\frac{p2^{p-2}}{6}\mathcal{M}_{p-2}(\mu)\mathcal{M}_{3}(V[\mu])t^{3}.$
Similarly, we obtain an estimate for $\biggl|\mathcal{M}_{p}\bigl(e^{V}_{t}(\mu)\bigr)-\mathcal{M}_{p}(\mu)\biggr|$.
Now, the first statement follows by induction using (3.84) in every step, and a subsequent application of Grönwall’s inequality.
To obtain the second estimate, we use (3.82) on $\mathbb{R}^{n}$ to obtain, in particular, the existence of a constant $\mathcal{H}^{\prime}$ depending only on $\mathcal{M}_{p}(\mu_{0})$, $T$, $R$ and $p$, and $\delta>0$, such that if the path–length $\mathfrak{p}(\mathcal{}\mathfrak{P})\leq\delta$, then, for all $t\in[0,T]$, $\bigl|\mathcal{M}_{p}(\mu^{\mathcal{}\mathfrak{P}}_{t})\bigr|\leq\mathcal{H}^{\prime}$. Combining this uniform upper bound of the moments with (3.84), the second statement readily follows by induction.
Finally, we obtain the estimates for an arbitrary manifold with bounded geometry by generalising the above proof analogously to the generalisation of the proof of Lemma 3.33 in Lemma 3.34. $\square$
Proceeding as in Lemma 3.34, we obtain uniform upper bounds for the difference of the pp–th and the qq–th moments of an AFAS.
@@
Let $(M,g)$ be of bounded geometry. Let $T>0$, $p\geq q\geq 2$, $\mu_{0}\in\mathcal{W}_{p}(M)$. Let $\mathcal{}\mathfrak{P}$ be a partition of $[0,T]$. Let $V$ be a probability vector–field. Assume that there exists $R>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $V[\mu]\in\mathcal{W}_{p+2}(\mathcal{X}^{2,\infty}(M))$ and $\mathcal{M}_{p+2}(V[\mu])\leq R^{p+2}$. Then, there exists $\delta>0$ and a constant $\mathcal{K}>0$ depending only on $\mathcal{M}_{p}(\mu_{0}),T,p,R$ and $M$, such that, if $\mathfrak{p}(\mathcal{}\mathfrak{P})\leq\delta$, for all $t\in[0,T]$,
$\biggl|\mathcal{M}_{p}(\mu^{\mathcal{}\mathfrak{P}}_{t})-\mathcal{M}_{q}(\mu^{\mathcal{}\mathfrak{P}}_{t})\biggr|\leq\biggl|\mathcal{M}_{p}(\mu_{0})-\mathcal{M}_{q}(\mu_{0})\biggr|+\mathcal{K}(p-q)t.$ (3.85)
@@
For $(M,g)=(\mathbb{R}^{n},g^{eucl})$, the result follows from a computation similar to the one in the proof of Lemma 3.33. Indeed,
@@
$\hskip 18.49988pt+\bigl(p\|x-x_{0}\|^{p-2}-q\|x-x_{0}\|^{q-2}\bigr)\bigl<x-x_{0},X(x)\bigr>t$
$\hskip 18.49988pt+\int_{0}^{t}(t-s)\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}\biggl(t\mapsto\bigl\|\rho^{X}_{t}(x)-x_{0}\bigr\|^{p}-\bigl\|\rho^{X}_{t}(x)-x_{0}\bigr\|^{q}\biggr)(s)\operatorname{d}s,$
@@
$\hskip 18.49988pt=\biggl(p(p-2)\bigl\|\rho^{X}_{s}(x)-x_{0}\bigr\|^{p-2}-q(q-2)\bigl\|\rho^{X}_{s}(x)-x_{0}\bigr\|^{q-2}\biggr)$
$\hskip 18.49988pt\cdot\biggl<\frac{\rho^{X}_{s}(x)-x_{0}}{\|\rho^{X}_{s}(x)-x_{0}\|},\frac{\operatorname{d}}{\operatorname{d}t}\biggl(t\mapsto\rho^{X}_{t}(x)\biggr)(s)\biggr>^{2}$
$\hskip 18.49988pt+\biggl(p\bigl\|\rho^{X}_{s}(x)-x_{0}\bigr\|^{p-2}-q\bigl\|\rho^{X}_{s}(x)-x_{0}\bigr\|^{q-2}\biggr)\biggl[\biggl\|\frac{\operatorname{d}}{\operatorname{d}t}\biggl(t\mapsto\rho^{X}_{t}(x)\biggr)(s)\biggr\|^{2}$
[…]
@@
$\hskip 18.49988pt\leq(p-q)\biggl|\log\bigl(\bigl\|\rho^{X}_{s}(x)-x_{0}\bigr\|\bigr)\biggr|$
$\hskip 18.49988pt\hskip 18.49988pt\cdot\max\biggl(\bigl\|\rho^{X}_{s}(x)-x_{0}\bigr\|^{p-2},\bigl\|\rho^{X}_{s}(x)-x_{0}\bigr\|^{q-2}\biggr).$
@@
$\biggl|\log\bigl(\bigl\|\rho^{X}_{s}(x)-x_{0}\bigr\|\bigr)\biggr|\bigl\|\rho^{X}_{s}(x)-x_{0}\bigr\|^{q-2}$ (3.89)
$\hskip 18.49988pt\leq\bigl\|\rho^{X}_{s}(x)-x_{0}\bigr\|^{q-2+\delta}+\bigl\|\rho^{X}_{s}(x)-x_{0}\bigr\|^{q-2-\delta}.$
@@
$\hskip 18.49988pt\leq\biggl|\mathcal{M}_{p}(\mu)-\mathcal{M}_{q}(\mu)\biggr|$
$\hskip 18.49988pt\qquad+2^{p-2}\bigl(8p^{2}R^{2}+6R^{2}p+3R^{2}p\bigr)\bigl(\mathcal{M}_{p}(\mu)+1\bigr)(p-q)t^{2}$
$\hskip 18.49988pt\qquad+\frac{2^{2(p-1)}\bigl(8p^{2}R^{2}+6R^{2}p+3R^{2}p\bigr)R^{2}}{6}(p-q)t^{3}.$
Now, the result follows by induction on the steps in $\mathcal{}\mathfrak{P}$ using (3.82).
Finally, we obtain the estimates for an arbitrary manifold with bounded geometry by generalising the above proof analogously to the generalisation of the proof of Lemma 3.33 in Lemma 3.34. $\square$
@@ 3.6.5 First Main Result: Generalised Central Limit Theorem—weak version
Let $T>0$ and $\mu_{0}\in\mathcal{W}_{p}(M)$. Let $\bigl(\mathcal{}\mathfrak{P}(\frac{1}{N})\bigr)_{N\in\mathbb{N}}$ be a family of partitions of $[0,T]$. Let $V$ be a uniformly continuous probability vector–field on $\mathcal{W}_{p}(M)$ for $p>2$. Assume that $V$ has uniformly bounded pp–th moments, e.g., there exists a constant $B>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $\mathcal{M}_{p}(V[\mu])<B$.
@@
    Assume that $(M,g)$ is of bounded geometry. Assume furthermore that there exists $R>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $V[\mu]\in\mathcal{W}_{p+2}(\mathcal{X}^{2,\infty}(M))$ and $\mathcal{M}_{p+2}(V[\mu])\leq R^{p+2}$. Then, there exists a subsequence $\biggl(\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N_{k}})}\biggr)_{k\in\mathbb{N}}$ of the Average Flow Approximation Series $\biggl(\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N})}\biggr)_{N\in\mathbb{N}}$ with initial value $\mu_{0}$ (see Definition 3.29) which uniformly converges with respect to the pp–Wasserstein distance to a $\frac{1}{p}$–Hölder–continuous smooth solution $\mu:[0,T]\to\mathcal{W}_{p}(M)$ of (MDEd) with initial value $\mu_{0}$.
@@
    Now, we only assume that there exists $q<p$ such that $V$ is continuous for the $\mathcal{W}_{q}$–distance. Then, there exists a subsequence $\biggl(\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N_{k}})}\biggr)_{k\in\mathbb{N}}$ of the Average Flow Approximation Series with initial value $\mu_{0}$ which converges to a smooth solution $\mu:[0,T]\to\mathcal{W}_{p}(M)$ of (MDEd) with initial value $\mu_{0}$.
@@
Let $p>2$. Take $N\in\mathbb{N}$, and let $(x_{i})_{i\in\llbracket 0,I_{N}\rrbracket}\coloneqq\mathcal{}\mathfrak{P}(\frac{1}{N})$ indexed in increasing order, To alleviate notation, we shall write $\mu^{N}\coloneqq\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N})}$. The proof consists of three main steps.
@@
Take $N\in\mathbb{N}$ and $l\in\llbracket 0,I_{N}-1\rrbracket$. Suppose $\frac{1}{2N}\leq 1$. Then for all $t,s\in\bigl[\frac{x_{l}+x_{l+1}}{2},x_{l+1}\bigr]$, we estimate $|t-s|\leq\sqrt{|t-s|}$. Note that since $V$ has pp–th moments bounded by $B>0$, it has in particular first and second moments bounded by $B+1$. Then, Lemma 3.31 yields that
$\mathcal{W}_{p}(\mu^{N}(t),\mu^{N}(s))$ $\leq 2\mathcal{M}_{1}\biggl(V\biggl[\mu^{N}\bigl(\frac{x_{l+1}+x_{l}}{2}\bigr)\biggr]\biggr)\sqrt{|t-s|}$ (3.91)
$\leq 2(B+1)\sqrt{|t-s|}$
Furthermore, by Lemma 3.32, for all $t,s\in\bigl[x_{l},\frac{x_{l}+x_{l+1}}{2}\bigr]$,
$\mathcal{W}_{p}(\mu^{N}(t),\mu^{N}(s))$ $\leq 2\sqrt{|t-s|}\biggl(\|\overline{V}[\mu^{N}(x_{l})]\|_{\infty}+\mathcal{M}_{p}(V[\mu^{N}(x_{l})])^{\frac{1}{p}}\biggr)$ (3.92)
$\leq 2\sqrt{|t-s|}\bigl(3(B+1)+B^{\frac{1}{p}}\bigr)$
Hence, there exists $\mathcal{B}>0$ such that, whenever $|t-s|\leq\frac{1}{2N}$, we have
@@
In the second and most technical step, we prove that the Average Flow Approximation Series from Definition 3.29 satisfies (MDEd) in the limit.
Take $N\in\mathbb{N}$, $s\in[0,T]$ and denote by $l^{N}(s)\in\llbracket 0,I_{N}-1\rrbracket$ the maximal integer satisfying $x_{l^{N}(s)}\leq s$ and $\varphi\in\mathcal{C}_{c}^{\infty}(M)$. Then,
@@
$\hskip 18.49988pt+\int_{M}\varphi\operatorname{d}(f^{V}_{x_{i+1}-x_{i}}(\mu^{N}(x_{i}))-\mu^{N}(x_{i}))$
@@
$\hskip 18.49988pt+\int_{M}\int_{\mathcal{X}^{2,\infty}(M)}\varphi\biggl(\rho^{\overline{X}[\mu^{N}(x_{i})]}_{\sqrt{x_{i+1}-x_{i}}}(x)\biggr)-\varphi(x)\operatorname{d}V[\mu^{N}(x_{i})](X)\operatorname{d}\mu^{N}(x_{i})(x).$
Now, note that for all $X\in\mathcal{X}^{2,\infty}(M)$, $d_{\infty}(\rho^{X}_{t},\operatorname{id})\leq t\|X\|_{\infty}$. Hence $\varphi\circ\rho^{X}_{t}$ is compactly supported. Furthermore, there exists a function $K:\mathcal{X}^{2,\infty}(M)\to\mathbb{R}$ bounded by $\overline{B}$ such that
@@
$\int_{\mathcal{X}^{2,\infty}(M)}\varphi(\rho^{\overline{X}[\mu^{N}(x_{i})]}_{\sqrt{x_{i+1}-x_{i}}}(x))-\varphi(x)\operatorname{d}V[\mu^{N}(x_{i})](X)$ (3.97)
$=\int_{\mathcal{X}^{2,\infty}(M)}\sqrt{x_{i+1}-x_{i}}\mathcal{L}_{\overline{X}[\mu^{N}(x_{i})]}\varphi(x)+\tfrac{1}{2}(x_{i+1}-x_{i})\mathcal{L}^{2}_{\overline{X}[\mu^{N}(x_{i})]}\varphi(x)$
$\hskip 18.49988pt+K(X)\|\varphi\|_{\mathcal{C}^{3}}(x_{i+1}-x_{i})^{\frac{\min(3,p)}{2}}\|{\overline{X}[\mu^{N}(x_{i})]}\|_{W^{2,\infty}}^{\min(3,p)}\operatorname{d}V[\mu^{N}(x_{i})](X)$
$=\int_{\mathcal{X}^{2,\infty}(M)}\sqrt{x_{i+1}-x_{i}}\mathcal{L}_{\overline{X}[\mu^{N}(x_{i})]}\varphi(x)$
$\hskip 18.49988pt\hskip 18.49988pt+\tfrac{1}{2}(x_{i+1}-x_{i})\mathcal{L}^{2}_{\overline{X}[\mu^{N}(x_{i})]}\varphi(x)\operatorname{d}V[\mu^{N}(x_{i})](X)$
$\hskip 18.49988pt+(x_{i+1}-x_{i})^{\frac{\min(3,p)}{2}}\!\!\int_{\mathcal{X}^{2,\infty}(M)}\!\!\!\!\!\!\!\!K(X)\|\varphi\|_{\mathcal{C}^{3}}\|{\overline{X}[\mu^{N}(x_{i})]}\|_{W^{2,\infty}}^{\min(3,p)}\operatorname{d}V[\mu^{N}(x_{i})](X).$
@@
$\int_{\mathcal{X}^{2,\infty}(M)}$ $\mathcal{L}_{\overline{X}[\mu^{N}(x_{i})]}\varphi(x)\operatorname{d}V[\mu^{N}(x_{i})](X)$ (3.98)
$=\int_{\mathcal{X}^{2,\infty}(M)}\operatorname{d}_{x}\varphi(\overline{X}[\mu^{N}(x_{i})](x))\operatorname{d}V[\mu^{N}(x_{i})](X)$
$=\int_{\mathcal{X}^{2,\infty}(M)}\operatorname{d}_{x}\varphi(v-\overline{V}[\mu^{N}(x_{i})](x))\operatorname{d}(\pi_{x})_{*}V[\mu^{N}(x_{i})](v)$
$=\operatorname{d}_{x}\varphi\biggl(\int_{\mathcal{X}^{2,\infty}(M)}v\operatorname{d}(\pi_{x})_{*}V[\mu^{N}(x_{i})](v)-\overline{V}[\mu^{N}(x_{i})](x)\biggr)=0,$